# Exciton fractional Chern insulators in moiré heterostructures

### v1

@@ Abstract
Moiré materials have emerged as a powerful platform for exploring exotic quantum phases. While recent experiments have unveiled fractional Chern insulators exhibiting the fractional quantum anomalous Hall effect based on electrons or holes, the exploration of analogous many-body states with bosonic constituents remains largely uncharted. In this work, we predict the emergence of bosonic fractional Chern insulators arising from long-lived excitons in a moiré superlattice formed by twisted bilayer WSe2 stacked on monolayer MoSe2. Performing exact diagonalization on the exciton flat Chern band present in this structure, we establish the existence of Abelian and non-Abelian phases at band filling $\frac{1}{2}$ and $1$, respectively, through multiple robust signatures including ground-state degeneracy, spectral flow, many-body Chern number, and particle-cut entanglement spectrum. The obtained energy gap of $\sim 10$ meV for the Abelian states suggests a remarkably high stability of this phase. Our findings not only introduce a highly tunable and experimentally accessible platform for investigating bosonic fractional Chern insulators but also open a new pathway for realizing non-Abelian anyons.
I Introduction
@@
Fractional Chern insulators (FCIs)—lattice analogs of the fractional quantum Hall effect that remain robust in the absence of a magnetic field—hold large potential for the study of fundamental quantum phenomena and for the development of novel quantum technologies [1, 2, 3]. Pioneering experiments [4, 5] and predictions [6, 7, 8] of FCIs in twisted van der Waals heterostructures, followed by recent realizations at absent magnetic field [9, 10, 11, 12], have established moiré materials as an accessible and versatile platform for exploring strongly-correlated topological phases. This breakthrough has stimulated abundant efforts in the pursuit of exotic phases beyond the conventional paradigm of Laughlin and hierarchy fractional quantum Hall states. In particular, predictions of Moore–Read [13, 14, 15, 16, 17, 18] and Read–Rezayi [19] phases hosting non-Abelian anyon excitations are especially promising for fault-tolerant topological quantum computing [20]. So far, however, research on moiré FCIs has focused only on the approach of doping the system with electrons or holes—resulting in correlated topological phases with fermionic constituents and leaving their bosonic counterpart largely unexplored.
@@
Excitons, i.e. Coulomb-bound pairs of conduction-band electrons and valence-band holes [21], are obvious candidates for realizing correlated bosonic phases in moiré materials [22, 23]. Concretely, interlayer excitons with charge-carriers located in different layers are particularly promising due to their long lifetime, which can reach hundreds of nanoseconds [24, 25] and even microseconds [26]. These species typically appear in semiconducting van der Waals heterostructures, where an optical excitation generating intralayer excitons (with electrons and holes in the same layer) is shortly followed by tunneling of either electrons or holes into another layer [27]. The resulting interlayer excitons possess a permanent dipole moment, which makes them highly tunable by an out-of-plane electric field [28, 29]. Moreover, repulsive dipolar interactions together with the presence of flat bands in moiré semiconductors lead to correlated exciton physics  [30, 31, 32, 33]. In addition, the large degree of control via twist angle, dielectric environment, and electric field can be exploited to achieve topological bands [34] hosting long-lived interlayer excitons [35]. Thus, moiré systems enable the realization of topological exciton flat bands, providing a promising route for the exploration of exciton physics with intertwined correlations and topology.
Figure 1: (a) Schematic illustration of twisted bilayer WSe2 on top of monolayer MoSe2. The interlayer excitons X1 and X2 are formed by an electron in MoSe2 and a hole in either of the two WSe2 layers. (b) Exciton band structure for $\Delta=-3.8$ meV and $\theta=1.95^{\circ}$, where the lowest band is flat and has a non-zero Chern number. The color represents the contribution $|\chi_{l\mathbf{Q}}|^{2}$ from each exciton species to the band. (c) Berry curvature $\Omega_{\mathbf{k}}A_{\text{BZ}}/2\pi$ and (d) Fubini-Study metric $\text{tr}[g_{\mathbf{k}}]A_{\text{BZ}}/2\pi$ in the moiré Brillouin zone for the flat band. $A_{\text{BZ}}$ is the Brillouin zone area.
@@
In this work, we unveil the many-body topological exciton phases emerging in a moiré heterostructure. First we show that, besides being topologically non-trivial, the lowest exciton band in twisted bilayer WSe2 stacked on monolayer MoSe2 can simultaneously be nearly flat and exhibit an almost ideal quantum geometry—posing this system as a strong candidate for realizing exciton FCI phases. Employing exact diagonalization, we show that the ground state at half filling with contact interactions is analogous to the bosonic Laughlin state in the lowest Landau level. Concretely, the state is characterized by a twofold degeneracy as well as an approximately zero (interaction) energy, and its nature is further confirmed by the many-body Chern number, spectral flow, and entanglement spectrum. Importantly, the Laughlin states remain robust and exhibit a large gap ($\sim 10$ meV) when replacing the idealized contact potential by realistic long-range interactions. Finally, the many-body calculations at filling $1$, again considering long-range interactions, provide compelling evidence for a stable bosonic version of the non-Abelian Moore–Read state. Overall, our findings predict the existence of exciton FCIs in an experimentally accessible moiré heterostructure and pave the way for the realization of Abelian and non-Abelian topological phases with bosonic constituents in moiré materials.
II Model
II.1 Topological exciton flat band
@@
We consider a van der Waals heterostructure consisting of a twisted WSe2 bilayer stacked on top of a single MoSe2 layer. An optical excitation and subsequent tunneling result in the formation of long-lived interlayer excitons composed of a hole in either of the two WSe2 layers and an electron in MoSe2, cf. Fig. 1(a). We assume spin-valley polarization of electrons and holes, which can be achieved by a circularly-polarized optical excitation if the valley decoherence time is sufficiently long and, in some cases, occurs spontaneously due to interactions. In order to describe moiré excitons in this structure, we follow the procedure introduced in Ref. [35].
First, we set up the exciton basis, |$\ket{X_{l\mathbf{Q}}}=\sum_{\mathbf{k}}\phi_{l\mathbf{k}}e^{\dagger}_{\mathbf{k}_{\text{e}}}h^{\dagger}_{l\mathbf{k}_{\text{h}}}\ket{0}$, which accounts for the binding of an MoSe2 electron to a WSe2 hole in the layer $l=1,2$. Details of the exciton basis, including the relation between the electron/hole momentum $\mathbf{k}_{\text{e/h}}$ and the relative and center-of-mass momenta $\mathbf{k}$ and $\mathbf{Q}$ can be found in Appendix A.1. The exciton’s wave function $\phi_{l\mathbf{k}}$ and binding energy $E^{\text{x}}_{l}$ are obtained by solving the Schrödinger equation for an electron–hole pair interacting via the interlayer 2D Coulomb potential. Here, we describe electrons and holes in an effective-mass approximation, where the exciton eigenvalue problem takes the form of the Wannier equation. In this case, $\phi_{l\mathbf{k}}$ is well described by the 1s wave function of 2D hydrogen, and variational minimization considering material-specific parameters yields the exciton binding energies $E^{\text{x}}_{1}=230\ \text{meV},E^{\text{x}}_{2}=140\ \text{meV}$ in close agreement with the exact solutions.
Next, we incorporate the effect of the moiré potential and tunneling that holes in the WSe2 layers experience, which is described by the moiré exciton Hamiltonian [35, 36]
$H_{\text{x,m}}=\sum_{ll^{\prime}\mathbf{Qq}}M_{ll^{\prime}\mathbf{Qq}}X^{\dagger}_{l\mathbf{Q+q}}X^{\phantom{\dagger}}_{l^{\prime}\mathbf{Q}},$ (1)
where mixing with higher exciton states (e.g. 2p) is expected to be weak and has therefore been disregarded. We note that this approach naturally accounts for the mixing between all moiré hole bands forming the exciton state. Here, $M_{ll\mathbf{Qq}}=E^{\text{x}}_{l\mathbf{Q}}\delta_{\mathbf{q},0}+\mathcal{F}_{ll,\alpha_{\text{e}}\mathbf{q}}U_{\mathbf{q}}$ contains the exciton dispersion $E^{\text{x}}_{l\mathbf{Q}}=E^{\text{gap}}_{l}+E^{\text{x}}_{l}+\hbar^{2}\mathbf{Q}^{2}/2M_{\text{x}}$ with the exciton mass $M_{\text{x}}$, the hole moiré potential $U_{\mathbf{q}}=U_{0}\sum_{n=1}^{6}\mathrm{e}^{i\sigma_{nl}\varphi}\delta_{\mathbf{q},\mathbf{g}_{n}}$ with $\sigma_{nl}=(-1)^{n+l-1}$, $\mathbf{g}_{n}=C^{n}_{6}\mathbf{g}_{0}$, $\mathbf{g}_{0}=4\pi/\sqrt{3}a_{\text{m}}\mathbf{\hat{x}}$, the moiré lattice constant $a_{\text{m}}$ and the exciton form factor $\mathcal{F}_{ll^{\prime},\mathbf{q}}=\sum_{\mathbf{k}}\phi^{*}_{l\mathbf{k}}\phi_{l^{\prime}\mathbf{k+q}}$. Tunneling between opposite layers is contained in $M_{ll^{\prime}\mathbf{Qq}}=\mathcal{F}_{ll^{\prime},\alpha_{\text{e}}\mathbf{q}}T_{\mathbf{q}}$ ($l\neq l^{\prime}$), where $T_{\mathbf{q}}=T_{0}\sum_{n=1}^{3}\delta_{\mathbf{q},{\bm{\kappa}}_{n}}$, ${\bm{\kappa}}_{n}=C^{n}_{3}(\mathbf{K}_{2}-\mathbf{K}_{1})$ and $\mathbf{K}_{l}$ is the K-point in layer $l$. The relevant parameters for holes in twisted bilayer WSe2 are $\varphi=128^{\circ}$, $U_{0}=9$ meV, and $T_{0}=18$ meV [35]. Diagonalizing Eq. (1) yields the band structure, which in this system is characterized by topological bands for a specific range of the twist angle and the offset between the two exciton species, $\Delta=E^{\text{x}}_{2}-E^{\text{x}}_{1}$. The latter can be experimentally controlled by an out-of-plane electric field.
@@
In this work, we consider the twist angle $\theta=1.95^{\circ}$ and the offset $\Delta=E^{\text{x}}_{2}-E^{\text{x}}_{1}=3.8$ meV, where the lowest band is flat and has a Chern number $C=1$, cf. Fig. 1(b). The non-zero Chern number can be understood in terms of a pseudospin, describing the superposition between the two interlayer exciton species |$\ket{X_{1\mathbf{Q}}}$ and |$\ket{X_{2\mathbf{Q}}}$, which wraps around the Bloch sphere once as $\mathbf{Q}$ traverses the moiré Brillouin zone. The varying pseudospin is encoded in the color of the flat band in Fig. 1(b), where the weight of the state |$\ket{X_{l\mathbf{Q}}}$, i.e. $|\chi_{l\mathbf{Q}}|^{2}$ with $\chi_{l\mathbf{Q}}$ being the moiré exciton eigenfunction, is shown. Note that $\chi_{l\mathbf{Q}}$ is a vector where each element accounts for the momentum $\mathbf{Q}+\mathbf{g}$ with $\mathbf{g}$ pointing to the outer moiré cells up to a certain cutoff. Interestingly, the quantum geometry of the band is nearly ideal [37, 38], i.e. $\mathrm{tr}[g_{\mathbf{k}}]\approx|\Omega_{\mathbf{k}}|$ where §$\textsection g_{\mathbf{k}}$ is the quantum (Fubini–Study) metric and $\Omega_{\mathbf{k}}$ is the Berry curvature, cf. Fig. 1 (c), (d). This property strongly suggests the emergence of zero-energy ground states at even-denominator filling of the band with pseudopotential interactions—in analogy to Laughlin states of bosons in Landau levels, albeit now in the absence of a magnetic field. A nearly ideal quantum geometry was also found to be the precursor of electron FCIs in twisted bilayer graphene [8, 39] despite the fact that fluctuations of the metric induce new competing states [40].
II.2 Many-body exciton Hamiltonian
The aim of this work is to unveil the phases emerging in a many-body system of excitons in the topological moiré flat band. To that end, we set up the many-body exciton Hamiltonian, $H_{\text{x}}=H_{\text{x,m}}+H_{\text{x-x}}$, where
describes the exciton-exciton interaction, $\rho_{l}(\mathbf{q})=\sum_{\mathbf{Q}}X^{\dagger}_{l\mathbf{Q+q}}X^{\phantom{\dagger}}_{l\mathbf{Q}}$ is the exciton density operator, :::absent:::: : denotes normal order, and
$V_{ll^{\prime}}(\mathbf{q})=$ $V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})+V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})-V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})-[V^{\text{e-h}}_{l^{\prime}l}(-\mathbf{q})]^{*}$
$V^{\text{e-e}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{h}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}$
$V^{\text{h-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{e}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|d|l-l^{\prime}|}$
$V^{\text{e-h}}_{ll^{\prime}}(\mathbf{q})=$ $F_{ll^{\prime}}(\alpha_{\text{e}}\mathbf{q},\alpha_{\text{h}}\mathbf{q})\frac{e_{0}^{2}}{2A\epsilon\epsilon_{0}|\mathbf{q}|}\mathrm{e}^{-|\mathbf{q}|dl}$ (3)
is the interaction potential, which contains individual contributions accounting for interactions between the charges in different excitons, and where $F_{ll^{\prime}}(\mathbf{q}_{1},\mathbf{q}_{2})=\mathcal{F}_{ll,\mathbf{q}_{1}}\mathcal{F}_{l^{\prime}l^{\prime},-\mathbf{q}_{2}}$ has been introduced. The derivation of the exciton–exciton interaction potential is outlined in Appendix A.2. Here, we have focused on the direct interaction, which at long distances behaves as the interaction between two dipoles with lengths $dl$ and $dl^{\prime}$, i.e. $V_{ll^{\prime}}(\mathbf{r})\sim(dl)(dl^{\prime})/|\mathbf{r}|^{3}$. Exchange interactions typically play a minor role in the context of interlayer excitons [41, 42, 43, 44].
Importantly, the exciton operators $X^{\dagger}_{\mathbf{Q}}$ and $X^{\phantom{\dagger}}_{\mathbf{Q}}$ obey bosonic commutation relations. Corrections to the bosonic character arising from the exciton’s fermionic substructure are typically expected in the regime where the distance between neighboring excitons is comparable to the size of an exciton [45], i.e. $n_{\text{x}}a_{\text{B}}^{2}\sim 1$ where $n_{\text{x}}$ is the exciton density and $a_{\text{B}}$ the exciton Bohr radius. In typical moiré systems where there is one exciton per moiré site, $n_{\text{x}}a_{\text{B}}^{2}\sim 0.01$ and thus the bosonic description is appropriate.
The many-body problem of interacting excitons is solved via exact diagonalization. Concretely, we project the exciton Hamiltonian $H_{\text{x}}$ into the topological flat band and we diagonalize $H_{\text{x}}$ in a finite-size system consisting of $N_{\text{x}}$ excitons in $N_{\text{s}}$ moiré sites such that the band filling is $\nu=N_{\text{x}}/N_{\text{s}}$. We distinguish exciton FCIs from competing orders by assessing the ground state degeneracy, many-body Chern number (defined in Appendix B.1), and quasi-hole excitations. The latter can be accessed via the particle-cut entanglement spectrum (PES) [46, 47], where the many-body system is partitioned into $N_{A}$ and $N_{B}=N_{\text{x}}-N_{A}$ particles. The set of eigenvalues $\{\xi\}$ of $-\mathrm{log}\rho_{A}$ consitute the PES, where $\rho_{A}=\text{tr}_{B}[\frac{1}{N_{d}}\sum_{i}^{N_{d}}|\Psi_{i}\rangle\langle\Psi_{i}|]$ is the reduced density matrix of the subsystem $A$, $|\Psi_{i}\rangle$ is the $i$-th ground state, and $N_{d}$ is the ground state degeneracy. Writing the many-body ground state as |$\ket{\Psi}=\sum_{j}\mathrm{exp}(-\xi_{j}/2)\ket{\Psi^{A}_{j}}\otimes\bra{\Psi^{B}_{j}}$ |, one can see that lower values of $\xi_{j}$ indicate a large weight of a specific subspace configuration |$\ket{\Psi^{A}_{j}}$—meaning that such configuration is allowed. Thus, the PES probes quasi-hole excitations that obey the generalized statistics of a specific quantum phase.
@@ II.2 Many-body exciton Hamiltonian
Figure 2: Laughlin states at half filling. (a) Many-body energy spectrum containing the 10 lowest energies for each momentum sector, (b) spectral flow, (c) particle-cut entanglement spectrum, and (d) many-body Berry curvature for the two ground states considering contact interactions without kinetic energy effects at $\nu=\frac{1}{2}$. The respective data considering the long-range interaction $V_{ll^{\prime}}(\mathbf{q})$ and the kinetic energy is shown in (e)-(h). The ground states in (a) and (e) are marked in red and have an average many-body Chern number $C_{\text{avg}}=1/2$. In the PES, the number of states below the first entanglement gap (denoted by the red solid line) is $1287$, matching the number of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states. The considered system has $N_{\text{s}}=18$ moiré sites and its spanning vector can be found in the Appendix B.2
III Abelian states at half filling
@@
Based on the criteria for how closely a moiré band can mimic a Landau Level—(i) nontrivial topology, (ii) flat dispersion, and (iii) ideal quantum geometry [37, 3, 38]—the exciton flat band shown in Fig. 1(b) appears to be an excellent candidate for realizing FCI phases analogous to the fractional quantum Hall effect of bosons. In this context, Laughlin states at half filling constitute a prototypical phase [48, 49]. They emerge as exact zero-energy states in a system of bosons interacting via a delta-function potential (i.e. contact interaction) in the lowest Landau level and their elementary excitations behave as Abelian anyons. We note that bosonic Laughlin states have also been proposed in a different setting, where the exciton-like bosons correspond to low-energy excitations of a fully-filled valley-polarized electron band [50, 51, 52]. In the following, we provide numerical evidence showing that the nearly ideal exciton Chern band indeed supports zero-energy Laughlin states for contact interactions. Subsequently, we will demonstrate that this FCI phase survives when considering the realistic long-range interactions introduced in Eq. (3).
@@
In order to strengthen the ideal conditions of the system and establish the emergence of Laughlin states in the exciton Chern band, we initially consider contact interactions $V_{ll^{\prime}}(\mathbf{q})=V_{0}$ and disregard the impact of the kinetic energy. With these assumptions, exact diagonalization yields two many-body ground states with approximately zero energy, cf. Fig. 2(a). Importantly, the ground states are separated from excited states by a large energy gap, which reflects the expected topological protection. The twofold degeneracy and the total momentum of each state are characteristic of Laughlin states and can be understood with the aid of Landau level physics in the thin-torus limit [53]. In particular, the exclusion principle that arises in the thin-torus limit for bosons in the lowest Landau level dictates that a zero-energy state contains, at most, one particle in two consecutive sites [54]. The two states that fulfill this principle and therefore constitute the two degenerate ground states are those with a Fock-space configuration $101010\cdots$ and its translational-invariant partner $010101\cdots$, whose total momentum matches that of our numerical ground states. Furthermore, upon threading a magnetic flux (corresponding to twisted boundary conditions), the ground states evolve into each other after one flux quantum and return to their original states after the insertion of two flux quanta (Fig. 2(b)), indicating that the two states are adiabatically connected and that the Hall conductivity is quantized to $\frac{1}{2}$. The latter aspect is further confirmed by a direct calculation of the many-body Chern number, which yields $C_{\text{avg}}=\frac{1}{2}$ for each ground state and arises from a homogeneous many-body Berry curvature, cf. Fig. 2(d).
Next, we calculate and analyze the PES, which reveals the nature of the elementary quasi-hole excitations of ground states. In particular, the states with lowest eigenvalues in the PES typically correspond to hole excitations that fulfill the generalized exclusion principle and therefore serve as a fingerprint for distinguishing a specific phase from competing orders. In Fig. 2(c), we display the PES where the subsystem $A$ is taken to consist of $N_{A}=4$ particles. Here, states with low eigenvalues are well separated from higher eigenvalues by a large entanglement gap. Importantly, the number of states below the gap exactly matches the analytical counting of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states. Thus, the many-body ground state degeneracy, spectral flow, many-body Chern number, and PES all taken together clearly confirm the emergence of a stable FCI phase of excitons in the half-filled moiré Chern band.

### v2

@@ Abstract
Moiré materials have emerged as a powerful platform for exploring exotic quantum phases. While recent experiments have unveiled fractional Chern insulators exhibiting the fractional quantum anomalous Hall effect based on electrons or holes, the exploration of analogous many-body states with bosonic constituents remains largely uncharted. In this work, we predict the emergence of bosonic fractional Chern insulators arising from long-lived excitons in a moiré superlattice formed by twisted bilayer WSe2 stacked on monolayer MoSe2. Performing exact diagonalization on the exciton flat Chern band present in this structure, we provide compelling evidence for the existence of Abelian and non-Abelian phases at band filling $\frac{1}{2}$ and $1$, respectively, through multiple robust signatures including ground-state degeneracy, spectral flow, many-body Chern number, and particle-cut entanglement spectrum. The obtained energy gap of $\sim 10$ meV for the Abelian states suggests a remarkably high stability of this phase, which persists for a relatively wide range of twist angles and vertical electric fields. Our findings establish the presence of robust bosonic fractional Chern insulators in highly tunable and experimentally accessible moiré heterostructures and unveil a promising pathway for realizing non-Abelian anyons.
@@
Fractional Chern insulators (FCIs)—lattice analogs of the fractional quantum Hall effect that remain robust in the absence of a magnetic field—hold large potential for the study of fundamental quantum phenomena and for the development of novel quantum technologies [1, 2, 3]. Pioneering experiments [4, 5] and predictions [6, 7, 8] of FCIs in twisted van der Waals heterostructures, followed by recent realizations at absent magnetic field [9, 10, 11, 12], have established moiré materials as an accessible and versatile platform for exploring strongly-correlated topological phases. This breakthrough has stimulated abundant efforts in the pursuit of exotic phases beyond the conventional paradigm of Laughlin and hierarchy fractional quantum Hall states. In particular, predictions of Moore–Read (MR) [13, 14, 15, 16, 17, 18] and Read–Rezayi [19] phases hosting non-Abelian anyon excitations are especially promising for fault-tolerant topological quantum computing [20]. So far, however, research on moiré FCIs has focused only on the approach of doping the system with electrons or holes—resulting in correlated topological phases with fermionic constituents and leaving their bosonic counterpart largely unexplored.
@@
Excitons, i.e. Coulomb-bound pairs of conduction-band electrons and valence-band holes [21], are obvious candidates for realizing correlated bosonic phases in moiré materials [22, 23]. Concretely, interlayer excitons with charge-carriers located in different layers are particularly promising due to their long lifetime, which can reach hundreds of nanoseconds [24, 25] and even microseconds [26]. These species typically appear in semiconducting van der Waals heterostructures, where an optical excitation generating intralayer excitons (with electrons and holes in the same layer) is shortly followed by tunneling of either electrons or holes into another layer [27]. The resulting interlayer excitons possess a permanent dipole moment, which makes them highly tunable by an out-of-plane electric field [28, 29]. Moreover, repulsive dipolar interactions together with the presence of flat bands in moiré semiconductors lead to correlated exciton physics  [30, 31, 32, 33]. In addition, the large degree of control via twist angle, dielectric environment, and electric field can be exploited to achieve topological bands [34]. In particular, long-lived interlayer excitons in a moiré heterostructure consisting of twisted bilayer WSe2 on monolayer MoSe2 (tWSe2–MoSe2), which exhibit topological flat bands, have been proposed as a promising route for the exploration of exciton physics with intertwined correlations and topology [35].
Crucially, however, the emergence of FCI topological order is not at all guaranteed in a topological flat band, and even less so in the case of non-Abelian states which usually require artificially adjusted interactions. For instance, in order to emulate the lowest Landau level and thus host Laughlin zero modes in presence of short-range pseudopotential interactions, a band must have an ideal quantum geometry (or be vortexable) in addition to being flat and topologically non-trivial [36, 37, 38]. Even such ideal bands host non-FCI states in some cases [39, 40, 41], and, all the more surprising, FCI order can emerge in trivial bands [42]. Thus, a topological flat band by itself is definitely not a proof of FCI order, and the presence or absence of topological order must be addressed with many-body calculations.
@@
In this work, we unveil the many-body topological exciton phases emerging in a moiré heterostructure. First we show that, besides being a nearly-flat Chern band, the lowest exciton band in tWSe2–MoSe2 exhibits an almost ideal quantum geometry—posing this system as a strong candidate for realizing exciton FCI phases. Employing exact diagonalization, we show that the ground state at half filling with contact interactions is analogous to the bosonic Laughlin state in the lowest Landau level. Concretely, the state is characterized by a twofold degeneracy as well as an approximately zero energy, and its nature is further confirmed by the many-body Chern number, spectral flow, and entanglement spectrum. Importantly, the Laughlin states remain robust and exhibit a large gap ($\sim 10$ meV) when replacing the idealized contact potential by realistic long-range interactions. Furthermore, the gap remains sizable across a wide range of experimentally accessible twist angles and electric fields, where the latter controls the energy offset between the two interlayer excitons (X1 and X2 in Fig. 1(a)). Finally, the many-body calculations at filling $1$ provide compelling evidence for a stable bosonic version of the non-Abelian MR state. Overall, our findings predict the existence of exciton FCIs in an experimentally accessible moiré heterostructure and pave the way for the realization of Abelian and non-Abelian topological phases with bosonic constituents in moiré materials.
@@
Nearly-ideal exciton Chern band—We consider the van der Waals heterostructure tWSe2–MoSe2, where an optical excitation and subsequent tunneling result in the formation of long-lived interlayer excitons composed of a hole in either of the two WSe2 layers and an electron in MoSe2, cf. Fig. 1(a). The formation and equilibration of these excitons occurs on an ultrafast sub-picosecond equilibration time [27]—many orders of magnitude smaller than the (up to microsecond [26]) lifetime. Interlayer excitons can also be obtained directly by electrical doping of the electron and hole layers [43] typically separated by an hBN spacer. We assume spin-valley polarization of electrons and holes, which can be achieved by a circularly-polarized optical excitation if the valley decoherence time is sufficiently long and, in some cases, occurs spontaneously due to interactions.
In order to describe moiré excitons in this structure, we consider the model introduced in Ref. [35] and outlined in the Supplemental Material (SM) [44, 45, 46]. In particular, we diagonalize the moiré exciton Hamiltonian [35, 47] $H_{\text{x,m}}$, which describes how the two relevant interlayer exciton states are affected and coupled by the moiré potential and tunneling that holes in the WSe2 layers experience. The lowest band has a Chern number $C=1$ for a specific range of the twist angle and the energy offset between the two exciton species, $\Delta=E^{\text{x}}_{2}-E^{\text{x}}_{1}$. The latter can be experimentally controlled by an out-of-plane electric field.
@@
Here, we consider the twist angle $\theta=1.95^{\circ}$ and the offset $\Delta=3.8$ meV where, besides possessing a finite Chern number $C=1$, the lowest band is flat, cf. Fig. 1(b). Interestingly, the quantum geometry of the band is nearly ideal [36, 37], i.e. $\mathrm{tr}[g_{\mathbf{k}}]\approx|\Omega_{\mathbf{k}}|$ where $\textsection g_{\mathbf{k}}$ is the quantum (Fubini–Study) metric [48] and $\Omega_{\mathbf{k}}$ is the Berry curvature, cf. Fig. 1 (c), (d). This property strongly suggests the emergence of zero-energy ground states at even-denominator filling of the band with pseudopotential interactions—in analogy to Laughlin states of bosons in Landau levels, albeit now in the absence of a magnetic field. A nearly ideal quantum geometry was also found to be the precursor of electron FCIs in twisted bilayer graphene [8, 49] despite the fact that fluctuations of the metric induce new competing states [50].
Figure 1: Nearly ideal exciton Chern band. (a) Schematic illustration of the considered tWSe2–MoSe2 structure. The interlayer excitons X1 and X2 are formed by an electron in MoSe2 and a hole in either of the two WSe2 layers. (b) Exciton band structure for $\Delta=-3.8$ meV and $\theta=1.95^{\circ}$, where the lowest band has a Chern number $C=1$. The color represents the contribution from each exciton species to the band. (c) Berry curvature $\Omega_{\mathbf{k}}A_{\text{BZ}}/2\pi$ and (d) Fubini-Study metric $\text{tr}[g_{\mathbf{k}}]A_{\text{BZ}}/2\pi$ of the flat band across the moiré Brillouin zone. $A_{\text{BZ}}$ is the Brillouin zone area.
@@
Figure 2: Laughlin states at half filling. (a) Many-body energy spectrum containing the 10 lowest energies for each momentum sector, (b) spectral flow, (c) particle-cut entanglement spectrum, and (d) many-body Berry curvature for the two ground states considering contact interactions without kinetic energy effects at $\nu=\frac{1}{2}$. The respective data considering the realistic long-range interaction and the kinetic energy is shown in (e)-(h). The ground states in (a) and (e) are marked in red and have an average many-body Chern number $C_{\text{avg}}=1/2$. In the PES, the number of states below the first entanglement gap (denoted by the red solid line) is $1287$, matching the number of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states. The considered system has $N_{\text{s}}=18$ moiré sites.
In order to elucidate whether the nearly ideal exciton Chern band indeed hosts FCI order, we will exactly diagonalize the many-body Hamiltonian $H_{\text{x}}=H_{\text{x,m}}+H_{\text{x-x}}$ projected into the lowest band. The interaction Hamiltonian, $H_{\text{x-x}}$, is described in the End Matter. We consider finite-sized systems consisting of $N_{\text{x}}$ excitons in $N_{\text{s}}$ moiré sites with the filling factor $\nu=N_{\text{x}}/N_{\text{s}}$. Importantly, the exciton operators $X^{(\dagger)}_{\mathbf{Q}}$ obey bosonic commutation relations. Corrections to the bosonic character arising from the exciton’s fermionic substructure are typically expected in the regime where the distance between neighboring excitons is comparable to the size of an exciton [51], i.e. $n_{\text{x}}a_{\text{B}}^{2}\sim 1$ where $n_{\text{x}}$ is the exciton density and $a_{\text{B}}$ the exciton Bohr radius. In typical moiré systems where there is one exciton per moiré site, $n_{\text{x}}a_{\text{B}}^{2}\sim 0.01$, the bosonic description is appropriate.
@@
Abelian states at half filling—Based on the criteria for how closely a moiré band can mimic a Landau Level—(i) nontrivial topology, (ii) flat dispersion, and (iii) ideal quantum geometry [36, 3, 37]—the exciton flat band shown in Fig. 1(b) appears to be an excellent candidate for realizing FCI phases analogous to the fractional quantum Hall effect of bosons. In this context, Laughlin states at half filling constitute a prototypical phase [52, 53]. Theoretically, they emerge as exact zero-energy states in a system of bosons with contact interactions in the lowest Landau level and their elementary excitations behave as Abelian anyons. Bosonic Laughlin states have also been proposed in a different setting, where the bosons correspond to particle-hole excitations out of a fully-filled valley-polarized electron band [54, 55]—however, these states require fine-tuned interactions in order to be stable, the systems do not offer a way of controlling the filling factor, and predictive numerical calculations are lacking. Additionally, composite-fermion mean-field theory based on a generic BCS-like Hamiltonian suggested the potential emergence of Laughlin states from the topological p-wave pairing of electrons and holes in an excitonic insulator [56]. Here, we consider long-lived interlayer excitons in the nearly ideal, yet realistic, exciton Chern band displayed in Fig. 1(b) and provide numerical evidence showing that it supports remarkably stable Laughlin states even for realistic long-range interactions.
@@
In order to strengthen the ideal conditions of the system and establish the emergence of Laughlin states in the exciton Chern band, we initially consider contact interactions and disregard the impact of the kinetic energy. With these assumptions, exact diagonalization yields two many-body ground states with approximately zero energy, cf. Fig. 2(a). Importantly, the ground states are separated from excited states by a large energy gap, which reflects the topological protection. The twofold degeneracy and the total momentum of each state are characteristic of Laughlin states and can be understood with the aid of Landau level physics in the thin-torus limit [57]. In particular, the thin-torus exclusion principle for $\nu=\frac{1}{2}$ Laughlin states dictates that a zero-energy state contains, at most, one particle in two consecutive sites [58]. The two states that fulfill this principle and therefore constitute the two degenerate ground states are those with a Fock-space configuration $101010\cdots$ and its translational-invariant partner $010101\cdots$, whose total momenta match those of our numerical ground states. Furthermore, upon threading a magnetic flux (corresponding to twisted boundary conditions), the ground states evolve into each other and return to their original states after the insertion of one and two flux quanta (Fig. 2(b)), respectively, indicating that the two states are adiabatically connected and that the Hall conductivity is quantized to $\frac{1}{2}$. The latter aspect is further confirmed by a direct calculation of the many-body Chern number, which yields $C_{\text{avg}}=\frac{1}{2}$ for each ground state and arises from a homogeneous many-body Berry curvature, cf. Fig. 2(d). Next, we calculate and analyze the particle-cut entanglement spectrum (PES), which reveals the nature of the elementary quasi-hole excitations in the ground states (see SM). The calculated PES exhibits a large entanglement gap, below which the number of states matches exactly the analytical counting of quasi-hole excitations in the $\nu=\frac{1}{2}$ Laughlin states (Fig. 2(c)). Thus, the many-body ground state degeneracy, spectral flow, many-body Chern number, and PES clearly confirm the emergence of a stable FCI phase of excitons in the half-filled moiré Chern band.